OpenStudy (camerondoherty):

Its asking a lot but... can someone guide me through a few of these questions and show me what to do?

OpenStudy (anonymous):

that's better

OpenStudy (camerondoherty):

Thanks c:

OpenStudy (anonymous):

^ ty organic Latte

OpenStudy (anonymous):

a polynomial of degree 3 can have at most 3 zeros it must have one zero it cannot have more than 3 because if you have say zeros at \(r_1,r_2,r_3\) then the polynomial will factor as \[p(x)=a(x-r_1)(x-r_2)(x-r_3)\]

OpenStudy (anonymous):

Is there a way for them both to be correct? Explain your answer. No a zero of a polynomial is a synonym for where it crosses the \(x\) axis if \(r\) is a zero then \(p(r)=0\) and so \((r,0)\) is on the graph

OpenStudy (camerondoherty):

So since it has a degree of 3 it crosses the x axis 3 times?

OpenStudy (anonymous):

i can cross once only like for example \(y=x^3\) only crosses once at \((0,0)\) http://www.wolframalpha.com/input/?i=y%3Dx^3

OpenStudy (camerondoherty):

Ohhh i see

OpenStudy (anonymous):

i can have two zeros, touch once and cross once http://www.wolframalpha.com/input/?i=y%3Dx^2%28x-2%29

OpenStudy (anonymous):

or it can cross three times http://www.wolframalpha.com/input/?i=y%3D%28x-2%29%28x%2B1%29%28x-4%29

OpenStudy (camerondoherty):

Thank You c:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

for part two i would pick \[g(x) = x^3 – x^2 – 4x + 4\] because it is easiest

OpenStudy (camerondoherty):

Then i graph and show the zeros right?

OpenStudy (anonymous):

a rational zero of this polynomial has to divide 4 it is pretty obvious that \(g(1)=1-1-4+4=0\)

OpenStudy (anonymous):

yeah you can graph it and also factor it here is the graph http://www.wolframalpha.com/input/?i=y%3D+x^3+%E2%80%93+x^2+%E2%80%93+4x+%2B+4

OpenStudy (camerondoherty):

So it has 3 zeros?

OpenStudy (anonymous):

you can see that the graph crosses the x axis at \(-2,1,2\) yes three zeros

OpenStudy (anonymous):

since those are the zeros that means \[g(-2)=0\\ g(1)=0\\ g(2)=0\] and it also means that it factors as \[g(x)=(x+2)(x-1)(x-2)\]

OpenStudy (camerondoherty):

And the y intercept is 5 right?

OpenStudy (anonymous):

the \(y\) intercept is at \(g(0)=4\) the constant, so no not \(5\)\ \((0,4)\) is on the graph

OpenStudy (camerondoherty):

Oh ok... So The Y intercept is 4 because that is where it intercepts the y axis, There are 3 zeros which are \[-2,1,2 \] But I don't get what it means by if it opens upwards or downwards :/

OpenStudy (anonymous):

i am not sure what that means either because it is not a quadratic

OpenStudy (camerondoherty):

Maybe because it shows up in the graph from the bottom?

OpenStudy (anonymous):

i would say it does neither and this is a mistake in the question you can see the graph from the wolfram link i sent

OpenStudy (anonymous):

i would say it does not open up or down, but goes from \(-\infty\) to \(+\infty\) that is more accurate

OpenStudy (camerondoherty):

Oh Ok Thanks c:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you good with this one 4) Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph.?

OpenStudy (camerondoherty):

Um... im not really sure... does it come with any equation?

OpenStudy (camerondoherty):

Oh you make an example?

OpenStudy (anonymous):

How do you find the x value of the vertex? the \(x\) values of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

you want to make up an example? if so make it easy like \(y=x^2+4x-2\) in which case the \(x\) coordinate of the vertex would be \[-\frac{4}{2\times 1}=-2\]

OpenStudy (camerondoherty):

Oh ok... do you simplify it?

OpenStudy (anonymous):

How do you find the y value of the vertex? That is easy enough, plug in the \(x\) values to find the \(y\) value in our example since the \(x\) coordinate of the vertex is \(-2\) the \(y\) coordinate is \[y=(-2)^2+4\times (-2)-2=4-8-2=-6\]

OpenStudy (anonymous):

lol you do not "simplify" it, you compute it

OpenStudy (camerondoherty):

well the bottom lol 2 times 1

OpenStudy (anonymous):

i picked an easy example so that it would be an integer, not a fraction

OpenStudy (camerondoherty):

Oh ok lol now i see XD silly me... but again it shows if it opens upwards or downwards...

OpenStudy (anonymous):

now this time that makes sense a parabola opens up or down lets graph the example i wrote http://www.wolframalpha.com/input/?i=y%3Dx^2%2B4x-2

OpenStudy (anonymous):

it is real real easy to tell if it opens up or opens down \[y=ax^2+bx+c\] opens up if \(a>0\) and opens down if \(a<0\) i.e. it opens up if the leading coefficient is positive and opens down if the leading coefficient is negative

OpenStudy (camerondoherty):

Ohhh ok

OpenStudy (anonymous):

How can he tell if the function has an even or odd degree? if it goes from \(-\infty\) to \(+\infty\) or vice versa, the degree is odd like \(y=x,y=x^3\) etc

OpenStudy (anonymous):

it it goes from \(+\infty \) to \(+\infty\) or \(-\infty\) to \(-\infty\) the degree is even like \(y=x^2\)

OpenStudy (anonymous):

Does this mean that he can determine the exact number of degrees the function has? hell no you could have a fifth degree polynomial that has 3 zeros just like a 3rd degree polynomial or a fourth degree polynomial that looks like a quadratic

OpenStudy (camerondoherty):

lol Thank You so very much... I really appreeciate it c:

OpenStudy (camerondoherty):

*apreciate

OpenStudy (anonymous):

yw hope it is more or less clear sorry about that stupid spam at the beginning i suspended them

OpenStudy (camerondoherty):

Lol Thanks... It was weird :s lol its really rude when people do that though...

OpenStudy (camerondoherty):

lol